# Self-Employment as a Signal: Career Concerns with Hidden Firm Performance

### v1

@@ 4 Equilibrium and characterization
This section characterizes equilibrium policies and wages under (i) a benchmark with naïve pricing and (ii) equilibrium pricing where firms infer from applications. We also establish the existence of an absorbing employment region and provide comparative statics in patience and risk aversion.
4.1 Benchmark (myopic) wage posting
Setup
In the benchmark (myopic) case, the posted wage depends only on the public state: [^13] Competition and free entry deliver zero expected profits. With a positive profit wedge $\pi$, the analysis goes through with $w^{N}=\mu-\pi$, shifting the cutoff monotonically.
@@ 4.1 Benchmark (myopic) wage posting
$\mathcal{A}^{N}\equiv\{\,(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{N}(\alpha,\beta)\,\}.$
Comparative statics at the benchmark
@@
Proof sketch. (i) follows from a monotone shift of $U_{E}$ in $w$; (ii) follows because the marginal value of a public signal (the option value embedded in $U_{S}$) scales with $\delta$ while $U_{E}$ holds the state fixed; (iii) is immediate from the comparison $u(\mu)$ vs. $\theta$ at the cutoff. Full proofs appear in Appendix A. ∎
Dynamics across states
As $(\alpha+\beta)$ grows, the one-step drift in beliefs from a public outcome, $|\mu^{\pm}-\mu|$, becomes smaller; this reduces the informational option value of self-employment and expands $\mathcal{A}^{N}$. Thus, late in a career with a long public track record, the incentives to keep producing public signals weaken.
4.2 Equilibrium wage posting and inference from applications
Endogenous applicant pools
With equilibrium firms, wages reflect selection into employment. Given a cutoff policy $\hat{\theta}(\alpha,\beta)$, the applicant pool at state $(\alpha,\beta)$ is $\Theta^{E}(\alpha,\beta)=[0,\hat{\theta}(\alpha,\beta)]$, and the competitive wage equals the pool mean talent under the Beta posterior truncated to this interval:
@@ 4.2 Equilibrium wage posting and inference from applications
For any $(\alpha,\beta)$ and $\delta\in(0,1)$, there exists a unique cutoff $\hat{\theta}^{S}(\alpha,\beta)\in[0,1]$ solving (10). The corresponding equilibrium wage is $w^{S}(\alpha,\beta;\hat{\theta}^{S})$ from (9). The employment region $\mathcal{A}^{S}\equiv\{(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{S}(\alpha,\beta)\}$ is absorbing.
@@
Naïve versus equilibrium
Equilibrium pricing internalizes adverse selection: $w^{S}(\alpha,\beta;\hat{\theta})\leq w^{N}(\alpha,\beta)$ with strict inequality whenever $\hat{\theta}<1$. Lower wages tighten the employment condition and shift the cutoff down.
@@ 4.3 Comparative statics and testable implications
Patience
Increasing $\delta$ raises the option value of generating public signals. In both pricing regimes, $\hat{\theta}(\alpha,\beta)$ is weakly decreasing in $\delta$. [^16] At the equilibrium regime this follows from the envelope property of $V$ and the fact that employment leaves the public state unchanged while self-employment stochastically moves it toward higher $\mu$ with probability $\theta$.
Risk aversion
Holding the wage schedule fixed, greater concavity of $u$ weakly increases $\hat{\theta}$ (employment becomes less attractive at a given wage because $u(w)$ falls). With equilibrium pricing, an additional (general-equilibrium) effect operates through composition: when fewer types apply, the applicant pool is stronger, which increases $w^{S}$ and partially offsets the direct effect. Overall, $\hat{\theta}^{S}$ remains weakly increasing in risk aversion under standard regularity (Appendix C).
Public-track-record length
As $(\alpha+\beta)$ increases, the one-step drift $|\mu^{\pm}-\mu|$ shrinks, reducing the informational benefit of an additional public signal. Hence $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$, and the absorbing employment region expands with career length.
Cross-sectional predictions
The model predicts: (i) stronger reputations (higher $\mu$) are associated with higher acceptance of employment offers when pricing is naïve, but with a muted pattern under equilibrium pricing; (ii) transitions from self-employment to opaque employment are more common after runs of low public performance; (iii) wage offers by equilibrium firms are lower, on average, than naïve benchmark wages at the same public state, reflecting adverse selection.
A note on quitting and off-path beliefs
@@
5 Three-period illustration
To visualize the mechanisms, consider a finite horizon with three decision dates $t=0,1,2$ (the last action occurs at $t=2$). The worker discounts with factor $\delta\in(0,1)$. For transparency and closed forms, suppose the prior at $t=0$ is uniform, $\text{Beta}(1,1)$, and the period utility over certain consumption $x\in[0,1]$ is CRRA on $[0,1]$, $u(x)=x^{\rho}$ with $\rho\in(0,1]$; [^17] At $\rho=1$ (risk neutrality) the period utility comparison collapses to means; the qualitative difference between benchmark and equilibrium posting still comes entirely from application-based inference. recall $u(1)=1,u(0)=0$, so $\mathbb{E}[u(y_{t})\mid S,\theta]=\theta$.
5.1 Last period ($t=2$)
At $t=2$ there is no continuation value. The worker compares $S$ (utility $\theta$) to $E$ (utility $u(w_{2})$), so the cutoff is $\hat{\theta}_{2}=u(w_{2})$.
Naïve pricing
Wages equal the public posterior mean at the current state:
$w^{N}_{2}(\alpha,\beta)=\mu(\alpha,\beta)=\frac{\alpha}{\alpha+\beta}.$
From the uniform prior, the three relevant states and wages are
$(\alpha,\beta)\in\{(1,1),(2,1),(1,2)\},\quad w^{N}_{2}\in\Big{\{}\tfrac{1}{2},\tfrac{2}{3},\tfrac{1}{3}\Big{\}}.$
Hence the last-period cutoffs under naïve pricing are
$\hat{\theta}^{N}_{2}(1,1)=(\tfrac{1}{2})^{\rho},\qquad\hat{\theta}^{N}_{2}(2,1)=(\tfrac{2}{3})^{\rho},\qquad\hat{\theta}^{N}_{2}(1,2)=(\tfrac{1}{3})^{\rho}.$
Equilibrium pricing
Firms infer from applications. With a cutoff policy $\hat{\theta}$, the last-period wage equals the truncated-posterior mean on $0,\hat{\theta}$]. For $\text{Beta}(\alpha,\beta)$ this mean is
$w^{S}_{2}(\alpha,\beta;\hat{\theta})\;=\;\frac{B(\hat{\theta};\,\alpha+1,\beta)}{B(\hat{\theta};\,\alpha,\beta)},$
with $B(\cdot)$ the incomplete beta function. The cutoff solves the fixed point
$\hat{\theta}^{S}_{2}(\alpha,\beta)=u\!\left(w^{S}_{2}(\alpha,\beta;\hat{\theta}^{S}_{2}(\alpha,\beta))\right).$
From the uniform prior:
$(\alpha,\beta)=(1,1):\quad$ $w^{S}_{2}(\hat{\theta})=\tfrac{1}{2}\hat{\theta}\ \Rightarrow\ \hat{\theta}^{S}_{2}(1,1)=\Big{(}\tfrac{\hat{\theta}}{2}\Big{)}^{\rho}\ \Rightarrow\ \hat{\theta}^{S}_{2}(1,1)=2^{-\rho/(1-\rho)}\quad(\rho<1),$
$(\alpha,\beta)=(2,1):\quad$ $w^{S}_{2}(\hat{\theta})=\tfrac{2}{3}\hat{\theta}\ \Rightarrow\ \hat{\theta}^{S}_{2}(2,1)=\Big{(}\tfrac{2}{3}\hat{\theta}\Big{)}^{\rho}\ \Rightarrow\ \hat{\theta}^{S}_{2}(2,1)=\left(\tfrac{2}{3}\right)^{\rho/(1-\rho)},$
$(\alpha,\beta)=(1,2):\quad$ $w^{S}_{2}(\hat{\theta})=\frac{\hat{\theta}-\frac{2}{3}\hat{\theta}^{2}}{\,2-\hat{\theta}\,}\ \Rightarrow\ \hat{\theta}^{S}_{2}(1,2)\text{ solves }\ \hat{\theta}=\Big{(}\frac{\hat{\theta}-\frac{2}{3}\hat{\theta}^{2}}{\,2-\hat{\theta}\,}\Big{)}^{\rho}.$
These solutions are interior for $\rho\in(0,1)$ and satisfy $\hat{\theta}^{S}_{2}(\alpha,\beta)\leq\hat{\theta}^{N}_{2}(\alpha,\beta)$, with equality only if $\hat{\theta}=1$ (no selection).
Numerical illustration
With $\rho=\tfrac{1}{2}$ (square-root) and, say, $\delta=0.95$ for the next subsection, the last-period cutoffs are:
$\hat{\theta}^{N}_{2}(1,1)\approx 0.707,\quad\hat{\theta}^{N}_{2}(2,1)\approx 0.817,\quad\hat{\theta}^{N}_{2}(1,2)\approx 0.577;$
$\hat{\theta}^{S}_{2}(1,1)=2^{-1}=0.5,\quad\hat{\theta}^{S}_{2}(2,1)=(2/3)^{1}=0.667,\quad\hat{\theta}^{S}_{2}(1,2)\approx 0.451.$
Thus sophistication lowers wages and cutoffs (relative to naïve), especially at low-belief states.
5.2 Middle period ($t=1$) and backward induction
Let $V_{2}(\theta;\alpha,\beta)=\max\{\theta,\ u(w_{2}(\alpha,\beta))\}$ denote the last-period value. At $t=1$ the worker compares
$U_{S}^{(1)}(\theta;\alpha,\beta)=\theta+\delta\Big{[}\,\theta\,V_{2}(\theta;\alpha+1,\beta)+(1-\theta)\,V_{2}(\theta;\alpha,\beta+1)\,\Big{]},$
$U_{E}^{(1)}(\theta;\alpha,\beta)=u(w_{1}(\alpha,\beta))+\delta\,V_{2}(\theta;\alpha,\beta),$
with $w_{1}=w^{N}$ or $w^{S}(\cdot;\hat{\theta}_{1})$ depending on the pricing regime. Strict single crossing in $\theta$ implies a unique middle-period cutoff $\hat{\theta}_{1}(\alpha,\beta)$.
Naïve pricing, uniform prior
At state $(1,1)$ we have $w^{N}_{1}=1/2$ and
$\begin{split}V_{2}(\theta;2,1)&=\max\{\theta,(2/3)^{\rho}\},\\ V_{2}(\theta;1,2)&=\max\{\theta,(1/3)^{\rho}\},\\ V_{2}(\theta;1,1)&=\max\{\theta,(1/2)^{\rho}\}.\end{split}$
The cutoff $\hat{\theta}^{N}_{1}(1,1)$ solves
$\hat{\theta}+\delta\Big{[}\hat{\theta}\,V_{2}(\hat{\theta};2,1)+(1-\hat{\theta})\,V_{2}(\hat{\theta};1,2)\Big{]}\;=\;(1/2)^{\rho}+\delta\,V_{2}(\hat{\theta};1,1).$
With $\rho=\tfrac{1}{2}$ and $\delta=0.95$, this yields [^18] In the relevant region $\hat{\theta}\in[0.577,0.707]$, $V_{2}(\hat{\theta};2,1)=(2/3)^{1/2}$, $V_{2}(\hat{\theta};1,2)=\hat{\theta}$, and $V_{2}(\hat{\theta};1,1)=(1/2)^{1/2}$, reducing the equation to a quadratic.
$\hat{\theta}^{N}_{1}(1,1)\approx 0.656.$
Equilibrium pricing, uniform prior
Now $w^{S}_{1}(1,1;\hat{\theta}_{1})=\mathbb{E}[\theta\mid\theta\leq\hat{\theta}_{1}]=\hat{\theta}_{1}/2$, so $u(w^{S}_{1})=\big{(}\hat{\theta}_{1}/2\big{)}^{\rho}$. The relevant last-period values use the state-specific equilibrium cutoffs from above:
$\begin{split}V_{2}^{S}(\theta;2,1)&=\max\{\theta,\ \hat{\theta}^{S}_{2}(2,1)\},\\ V_{2}^{S}(\theta;1,2)&=\max\{\theta,\ \hat{\theta}^{S}_{2}(1,2)\},\\ V_{2}^{S}(\theta;1,1)&=\max\{\theta,\ \hat{\theta}^{S}_{2}(1,1)\}.\end{split}$
At $(1,1)$ with $\rho=\tfrac{1}{2}$ and $\delta=0.95$, the fixed-point indifference equation becomes
$\hat{\theta}+\delta\Big{[}\hat{\theta}\,\max\{\hat{\theta},\,2/3\}+(1-\hat{\theta})\,\max\{\hat{\theta},\,0.451\}\Big{]}=\sqrt{\hat{\theta}/2}\;+\;\delta\,\max\{\hat{\theta},\,0.5\}.$
The unique solution is
$\hat{\theta}^{S}_{1}(1,1)\approx 0.42,$
substantially below $\hat{\theta}^{N}_{1}(1,1)\approx 0.656$, reflecting the adverse-selection discount in equilibrium wages.
Takeaways
Two features stand out. First, absorbing hidden employment: once a worker with $\theta\leq\hat{\theta}_{t}(\alpha,\beta)$ accepts $E$, the public state freezes and the inequality repeats, so employment is absorbing at that state. Second, pricing sophistication compresses wages and tightens self-selection: cutoffs fall ($\hat{\theta}^{S}\leq\hat{\theta}^{N}$), so a smaller set accepts employment early; mid-types instead generate public signals via $S$.
With $u(x)=\sqrt{x}$ and a uniform prior, equilibrium wage posting compresses the success–failure gap at $t=1$: benchmark wages are $2/3$ (after success) vs. $1/3$ (after failure), while equilibrium wages are about $0.407$ vs. $0.177$ under our calibration—a substantial winner’s-curse correction.
Acceptance-set evolution (equilibrium posting).
Under equilibrium wage posting, the applicant pool each date is $\Theta^{E}_{t}(\alpha,\beta)=[0,\hat{\theta}_{t}(\alpha,\beta)]$. Starting from $(1,1)$ at $t=1$, a public success (failure) moves the state to $(2,1)$ (resp. $(1,2)$), which induces distinct acceptance sets $0,\hat{\theta}_{1}(2,1)]$ and $0,\hat{\theta}_{1}(1,2)]$ and thus distinct wages $w^{E}_{1}(2,1)=\mathbb{E}[\theta\mid\theta\leq\hat{\theta}_{1}(2,1)]$ and $w^{E}_{1}(1,2)=\mathbb{E}[\theta\mid\theta\leq\hat{\theta}_{1}(1,2)]$. The same logic applies at $t=2$ for $(\alpha,\beta)\in\{(1,1),(2,1),(1,2)\}$ with $\hat{\theta}_{2}(\alpha,\beta)$ solving the corresponding fixed point. This makes explicit how acceptance sets and wages evolve along each branch of the public history.
@@ 6 Empirical implications
Increasing $\delta$ expands the option value of public signals. In both pricing regimes, the cutoff $\hat{\theta}(\alpha,\beta)$ is weakly decreasing in $\delta$: more patient workers are more willing to forgo insurance to refresh reputation.
@@
Holding the wage schedule fixed, greater concavity of $u$ (on $[0,1]$ with $u(0)=0,u(1)=1$) raises the utility of a given certain wage relative to a risky Bernoulli payoff with the same mean. Hence $\hat{\theta}(\alpha,\beta)$ is weakly increasing in risk aversion: more types accept employment at a given public state. Under equilibrium pricing, a composition effect operates (fewer applicants strengthen the pool and raise $w^{S}$), which reinforces the shift.
@@
As $(\alpha+\beta)$ grows, a single public outcome moves $\mu$ by less ($|\mu^{\pm}-\mu|$ shrinks), reducing the informational benefit of $S$. Thus $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$; absorbing hidden employment is more prevalent late in careers.
@@
Occupations with portable public histories (open-source, contests, platforms) should exhibit (i) higher early self-employment/portfolio-building by stronger types, (ii) sharper wage gains after publicly observed successes, and (iii) more frequent switches back to opaque employment following poor public realizations.
Application-based inference
With equilibrium pricing, application choices are informative. Relative to naïve environments (e.g., where HR ignores selection), we predict (i) lower average wages for otherwise identical public states, (ii) fewer mid-type applicants, and (iii) stronger sorting on outside reputations.
@@
The model yields simple tests: (i) event studies of wage paths after public successes/failures for workers who remain outside versus those who enter opaque employment; (ii) cross-occupation comparisons of the share of opaque spells by public-reputation quantiles; (iii) within-platform evidence that mid-quantile workers are more likely to keep generating public signals when employers are more “equilibrium” (e.g., settings with better screening or more experienced recruiters).

### v3

@@ 4 Equilibrium and characterization
This section characterizes equilibrium policies and wages under (i) a benchmark with naïve pricing and (ii) sophisticated pricing where firms infer from applications. We also establish the existence of an absorbing employment region and provide comparative statics in patience and risk aversion.
4.1 Myopic wage posting
In the benchmark case, the posted wage depends only on the public state: [^14] Competition and free entry deliver zero expected profits. With a positive profit wedge $\pi$, the analysis goes through with $w^{N}=\mu-\pi$, shifting the cutoff monotonically.
@@ 4.1 Myopic wage posting
$\mathcal{A}^{N}\equiv\left\{\,(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{N}(\alpha,\beta)\,\right\}.$
@@
Proof sketch. (i) follows from a monotone shift of $U_{E}$ in $w$; (ii) follows because the marginal value of a public signal (the option value embedded in $U_{S}$) scales with $\delta$ while $U_{E}$ holds the state fixed; (iii) is immediate from the comparison $u(\mu)$ vs. $\theta$ at the cutoff. ∎
As $(\alpha+\beta)$ grows, the length of the worker’s public track record increases and the posterior about talent becomes more precise. A single additional success or failure then moves beliefs by less than at earlier states. Holding the current mean belief $\mu(\alpha,\beta)$ fixed, the one-period option value of an extra public signal is therefore larger when the public record is short and smaller when it is long. Combined with risk aversion, this makes continued self-employment most attractive when the worker has accumulated relatively few public signals (the reputational stakes per signal are high), and makes hidden employment comparatively more attractive once a long track record is already in place. In other words, moving across states in the direction of larger $(\alpha+\beta)$ shifts more types into the absorbing hidden-employment region: at otherwise similar belief states, workers with long public histories are more willing to accept the firm’s insurance and stop generating signals than workers with very short histories.
4.2 Sophisticated wage posting and selection into employment
With sophisticated firms, wages reflect selection into employment: given a cutoff policy $\hat{\theta}(\alpha,\beta)$, the applicant pool at state $(\alpha,\beta)$ is $\Theta^{E}(\alpha,\beta)=[0,\hat{\theta}(\alpha,\beta)]$, and the competitive wage equals the pool mean talent under the $\mathrm{Beta}(\alpha,\beta)$ posterior truncated to this interval:
@@ 4.2 Sophisticated wage posting and selection into employment
For any $(\alpha,\beta)$ and $\delta\in(0,1)$, there exists a unique cutoff $\hat{\theta}^{S}(\alpha,\beta)\in[0,1]$ solving (10). The corresponding sophisticated wage is $w^{S}(\alpha,\beta;\hat{\theta}^{S})$ from (9). The employment region $\mathcal{A}^{S}\equiv\{(\alpha,\beta,\theta):\theta\leq\hat{\theta}^{S}(\alpha,\beta)\}$ is absorbing.
@@
Sophisticated pricing internalizes adverse selection: $w^{S}(\alpha,\beta;\hat{\theta})\leq w^{N}(\alpha,\beta)$ with strict inequality whenever $\hat{\theta}<1$. Lower wages tighten the employment condition and shift the cutoff down.
@@ 4.3 Comparative statics and testable implications
Increasing $\delta$ raises the option value of generating public signals. In both pricing regimes, $\hat{\theta}(\alpha,\beta)$ is weakly decreasing in $\delta$. [^17] At the sophisticated regime this follows from the envelope property of $V$ and the fact that employment leaves the public state unchanged while self-employment stochastically moves it toward higher $\mu$ with probability $\theta$.
Holding the wage schedule fixed, greater concavity of $u$ weakly increases $\hat{\theta}$ (employment becomes less attractive at a given wage because $u(w)$ falls). With sophisticated pricing, an additional (general-equilibrium) effect operates through composition: when fewer types apply, the applicant pool is stronger, which increases $w^{S}$ and partially offsets the direct effect. Overall, $\hat{\theta}^{S}$ remains weakly increasing in risk aversion under standard regularity.
As $(\alpha+\beta)$ increases, the one-step drift $|\mu^{\pm}-\mu|$ shrinks, reducing the informational benefit of an additional public signal. Hence $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$, and the absorbing employment region expands as the public track record becomes longer.
The cutoff $\hat{\theta}(\alpha,\beta)$ links cross-sectional variation in self-employment, wages, and information production. At a given public state $(\alpha,\beta)$, only types with $\theta>\hat{\theta}(\alpha,\beta)$ choose self-employment and continue to generate public signals; lower types choose hidden employment and stop. This implies that occupations or markets in which self-employment or public projects are more prevalent are precisely those in which the cutoff is relatively low and many types reveal themselves, whereas markets with high cutoffs feature both more opaque employment and less public information about talent. The distinction between naïve and sophisticated firms amplifies these patterns. Naïve firms post the public-mean wage $\mu(\alpha,\beta)$ and therefore pay the same wage to all workers with the same public record, regardless of how many low types would be willing to hide. Sophisticated firms anticipate selection into opaque jobs and post a lower wage in states where low types are more likely to accept employment, compressing wages but inducing more self-employment and more public signal production. Cross-sectionally, we therefore expect (i) more self-employment and richer public track records in markets where firms are sophisticated and performance technologies are transparent, and (ii) higher average wages but less public information in markets with naïve pricing and opaque technologies at the same public state.
@@
Taken together, these comparative statics suggest a simple answer to the question of which labor markets generate the most public information about talent. Markets with more patient and less risk-averse workers, with more transparent performance technologies, and with sophisticated rather than naïve wage setting all have smaller absorbing hidden-employment regions. By Proposition 3, higher patience and better insurance reduce the cutoff $\hat{\theta}(\alpha,\beta)$, while greater risk aversion and a lower reward for public output increase it; by Proposition 5, sophisticated firms choose lower wages than naïve firms at a given public state and thus induce more workers to remain in public production. In such markets, more types experiment with self-employment or public projects, generating richer public track records. Conversely, markets with impatient, highly risk-averse workers, opaque production technologies, and naïve wage setting have larger absorbing regions and slower public learning.
@@ 5 Empirical implications
The discount factor $\delta$ scales the value of future wage improvements that come from generating public signals. At a given public state $(\alpha,\beta)$, let $\Delta(\theta;\alpha,\beta,\delta)$ denote the expected discounted utility difference between choosing $S$ and $E$ for type $\theta$.
By construction, $\Delta$ can be written as a current-period payoff difference plus $\delta$ times the difference in continuation values. The current-period term does not depend on $\delta$, while the continuation term is strictly increasing in both $\theta$ and $\delta$: higher types benefit more from the option value of moving the public belief toward higher $\mu(\alpha,\beta)$, and this option value matters more when the worker is patient. As a result, $\Delta$ has increasing differences in $(\theta,\delta)$, and the unique cutoff $\hat{\theta}(\alpha,\beta;\delta)$ solving $\Delta(\hat{\theta};\alpha,\beta,\delta)=0$ is weakly decreasing in $\delta$.
More patient workers are therefore more willing to forgo the insurance of $E$ in order to refresh their reputation, and absorbing hidden employment is more prevalent when workers are impatient or face short effective horizons.
@@
Risk aversion makes the certain wage relatively more attractive than the risky Bernoulli payoff from $S$. To see this, consider a one-parameter family of utilities $\{u_{\rho}\}_{\rho}$ on $[0,1]$ with $u_{\rho}(0)=0$ and $u_{\rho}(1)=1$, where a higher $\rho$ indexes greater concavity in the sense of Pratt. At a fixed state $(\alpha,\beta)$ and wage schedule $w(\alpha,\beta)$, the current-period payoff under $E$ is the sure amount $u_{\rho}(w)$, whereas under $S$ it is the expected utility of a Bernoulli payoff with mean $\theta$.
Standard comparative statics of risk aversion imply that, holding the mean fixed, the certainty equivalent of the risky payoff falls as $\rho$ increases. Thus the gain from choosing $S$ over $E$ is pointwise decreasing in $\rho$, and the cutoff $\hat{\theta}(\alpha,\beta;\rho)$ that solves $\Delta(\hat{\theta};\alpha,\beta,\rho)=0$ is weakly increasing in $\rho$. More risk-averse workers are therefore more willing to accept hidden employment at a given public state.
Under sophisticated wage setting, a composition effect reinforces this comparative static: when more low- and mid-types enter $E$, the remaining pool of applicants is stronger, raising the sophisticated wage $w^{S}(\alpha,\beta)$ and further tilting the trade-off toward taking the firm job for marginal types.
@@
The length of the public record affects how informative a marginal success or failure is. With a Beta posterior, the public mean is $\mu(\alpha,\beta)=\alpha/(\alpha+\beta)$.
After a success or failure under $S$, the mean updates to
$\begin{split}\mu_{+}(\alpha,\beta)&=\frac{\alpha+1}{\alpha+\beta+1},\\ \mu_{-}(\alpha,\beta)&=\frac{\alpha}{\alpha+\beta+1},\end{split}$
so that
$\begin{split}\mu_{+}(\alpha,\beta)-\mu(\alpha,\beta)&=\frac{\beta}{(\alpha+\beta)(\alpha+\beta+1)},\\ \mu(\alpha,\beta)-\mu_{-}(\alpha,\beta)&=\frac{\alpha}{(\alpha+\beta)(\alpha+\beta+1)}.\end{split}$
For a given mean $\mu$, both expressions shrink like $1/(\alpha+\beta+1)$ as the effective sample size $(\alpha+\beta)$ grows. The option value of an extra public signal therefore declines with the length of the public track record.
Holding wages fixed, $\Delta(\theta;\alpha,\beta)$ shifts down as $(\alpha+\beta)$ increases, so the cutoff $\hat{\theta}(\alpha,\beta)$ is weakly increasing in $(\alpha+\beta)$. In steady state, absorbing hidden employment is thus more prevalent in parts of the labor market where typical workers already have long public histories, while early-career workers with short records have stronger incentives to continue producing public signals.
@@
These comparative statics carry clear cross-occupation implications. In occupations with portable public histories—open-source software, contests, online freelancing platforms, or fields with visible publication and citation records—workers can more easily accumulate long public track records.
Our model predicts (i) a higher prevalence of self-employment or public projects among workers with strong reputations in such occupations, because a lower cutoff $\hat{\theta}(\alpha,\beta)$ makes public production attractive for many types; (ii) sharper wage increases after publicly observed successes and steeper penalties after public failures, since wages respond more to public signals when they are informative; and (iii) more frequent transitions into opaque employment following runs of poor public realizations, especially for mid-ability workers whose reputations are pushed toward the absorbing hidden-employment region.
These patterns align with evidence that public ratings and track records on digital platforms and innovation ecosystems have first-order effects on hiring and earnings, and that workers respond by reallocating effort toward more visible tasks when the returns to transparency are high. [^18] See, for example, empirical work on online labor markets and open-source contributions that documents how public performance measures shape demand, wages, and workers’ portfolio choices.
Pricing sophistication and wage dispersion
Sophisticated wage setting makes the adverse selection in applicant pools explicit. In the naïve benchmark, wages equal the public posterior mean $w^{N}(\alpha,\beta)=\mu(\alpha,\beta)$ regardless of who is expected to apply.
Under sophisticated pricing, wages equal the mean talent of the implied applicant pool $0,\hat{\theta}(\alpha,\beta)]$, which is strictly below $\mu(\alpha,\beta)$ whenever $\hat{\theta}(\alpha,\beta)<1$. Thus at any given public state, sophisticated wages $w^{S}(\alpha,\beta)$ are lower on average than naïve benchmark wages, reflecting a familiar adverse-selection discount. This compresses wage dispersion after public successes and failures: when a positive public shock moves the state from $(\alpha,\beta)$ to $(\alpha+1,\beta)$, naïve wages jump by $\mu_{+}-\mu$, whereas sophisticated wages move by the change in the truncated-posterior mean.
Because sophisticated wages already incorporate who is expected to apply, the incremental impact of a public success on wages is smaller. Relative to naïve environments in which HR ignores selection into opaque jobs, sophisticated pricing therefore implies (i) lower average wages for otherwise identical public states, (ii) fewer mid-type applicants for opaque jobs (some mid-ability workers prefer to keep generating public signals to avoid the adverse-selection discount), and (iii) stronger sorting of workers into public versus hidden production based on outside reputations.
@@
The model suggests a set of empirical probes that map directly into the comparative statics above.
First, event-study designs can compare wage paths after public successes and failures for workers who remain in self-employment versus those who move into opaque employment, conditioning on the public state $(\alpha,\beta)$ at the time of the shock. The theory predicts larger wage responses and slower entry into opaque jobs after positive shocks in transparent, naïve-pricing environments, and more muted wage dispersion when pricing is sophisticated.
Second, cross-occupation comparisons can relate the share and timing of opaque spells to measures of transparency and public-track-record length: occupations with more portable public histories should exhibit earlier switching into hidden employment for weaker reputations and stronger persistence in opaque spells once workers enter the absorbing region.
Third, within-platform variation in “pricing sophistication”—for instance, across employers with different screening technologies or across platforms with more versus less curated rating systems—can be used to test whether mid-quantile workers are more likely to keep generating public signals when they anticipate stronger adverse-selection discounts in opaque jobs.
Together, these tests speak to how selection into self-employment versus hidden jobs and the sophistication of wage setting jointly determine both wage dynamics and the amount of public information that labor markets generate about talent.